OpenStudy (anonymous):

The finite sequence whose general term is an = 0.17n2 - 1.02n + 6.67 where n = 1, 2, 3, ..., 9 models the total operating costs, in millions of dollars, for a company from 1991 through 1999. Find (I will post pic)

OpenStudy (anonymous):

Answer choices A. $21.58 million B. $27.4 million C. $23.28 million D. $29.1 million

jimthompson5910 (jim_thompson5910):

find a1, a2, a3, a4, a5 Then add them up

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (anonymous):

how do you find a1 a2...

jimthompson5910 (jim_thompson5910):

To find a1, plug in n = 1 and evaluate an = 0.17n^2 - 1.02n + 6.67 a1 = 0.17(1)^2 - 1.02(1) + 6.67 a1 = 0.17(1) - 1.02(1) + 6.67 a1 = 0.17 - 1.02 + 6.67 a1 = 5.82

jimthompson5910 (jim_thompson5910):

Do the same thing, but using n = 2, to find a2

OpenStudy (anonymous):

for a2 i got 5.31 a3=5.14 a4=5.31 and a5=5.82 if you add a1 to a5 together you get 27.4

OpenStudy (anonymous):

so B?

jimthompson5910 (jim_thompson5910):

that is 100% correct

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

thnx!

jimthompson5910 (jim_thompson5910):

np